OpenStudy (anonymous):

Ramya paid 12.50 for a shirt and cap. The shirt was 4 times the cost of the cap. How much was the cap?

OpenStudy (across):

\[12.5=s+c\]\[s=4c\]\[12.5=4c+c\]\[12.5=5c\]\[c=2.5\]

OpenStudy (anonymous):

Solve the following for s and c. {s + c = 12.5, s = 4 c} s=10, c=2.5

OpenStudy (anonymous):

x=cost of cap so: 4x=cost of shirt so cost of cap and shirt equals price paid or: x+4x=$12.50 5x=$12.50 x=$2.50 so cap is $2.50 and shirt is 4 times that amount or4($2.50)=$10